OpenStudy (anonymous):

@johnweldon1993 I need help factoring again > . <

OpenStudy (johnweldon1993):

lol yes? :)

OpenStudy (anonymous):

One sec, writing em on sketchtoy. I have two problems :P

OpenStudy (johnweldon1993):

haha okay :P

OpenStudy (anonymous):

http://www.sketchtoy.com/40506599

OpenStudy (johnweldon1993):

Hang on ...taking time to load for me...

OpenStudy (johnweldon1993):

Okay...well that one is fairly simple... what are the factors of 15? and what are the factors of 2?

OpenStudy (johnweldon1993):

ACTUALLY!!! want to do the quadratic formula??? :) I know you asked me about it earlier...

OpenStudy (anonymous):

Let's do regular first, then quadratic. ^_^

OpenStudy (anonymous):

Factors of 15 are 3 and 5 and factors of 2 are 1 and 2

OpenStudy (johnweldon1993):

well ...the full factors are 15 = 1 , 3 , 5 , 15 ...but yes you chose the only 2 important ones here lol and yes 2 = 1, 2 so we know we need 2x² not just x² ...so there will not just be (x 3)(x 5) instead we will have (2x + 3)(x + 5) *we want positives because we have both positive numbers....so that's it...that's your factorization

OpenStudy (johnweldon1993):

Do you want to solve for the zeros?

OpenStudy (anonymous):

What do you mean, the zeros?

OpenStudy (johnweldon1993):

Here...the quadratic formula solves for the zeros of the function...so I'm sure that's what you wanted

OpenStudy (anonymous):

>.< uh ok

OpenStudy (johnweldon1993):

So we have (2x + 3)(x + 5) = 0 right? we can break that up 2x + 3 = 0 x + 5 = 0 solve for x in each equation

OpenStudy (johnweldon1993):

Lol you seem to be very confused based on that face lol....let me start over... you have 2x² + 13x + 15 = 0 to start with right?

OpenStudy (anonymous):

Yes

OpenStudy (johnweldon1993):

Okay...well see how it is set to = 0? that means we want to find out what values of 'x' make that equation = 0 okay? This is done by factoring! *this is what I meant when I said find the zeros

OpenStudy (johnweldon1993):

So....We factored....and now we are able to figure out exactly what values of 'x' make that equation = 0 (2x + 3)(x + 5) = 0 we know that with either of the parenthesis...we only need 1 to = 0 in order for everything to = 0 right? (0)(x + 5) = 0 because anything times 0 = 0 right? so we can split this into 2 equations 2x + 3 = 0 and x + 5 = 0 can you solve for 'x' in each of those equations?

OpenStudy (anonymous):

Both x's would be 0, right? > _ <

OpenStudy (johnweldon1993):

Not quite... 2x + 3 = 0 if 'x' = 0 then you would have 2(0) + 3 - 0 0 + 3 = 0 3 = 0 is this correct?

OpenStudy (anonymous):

Ohhh, you have to have it equal 1... right?

OpenStudy (johnweldon1993):

Here I'll do this one for you 2x + 3 = 0 you want to subtract 3 from both sides to get 2x by itself 2x + 3 = 0 -3 -3 2x = -3 divide both sides by 2 to completely isolate x... 2x = -3 -------- 2 2 x = -3/2 This is one "zero" meaning...if you plugged in (-3/2) into your original equation...you would get 0 which is what you want... so now x + 5 = 0 what does x = ?

OpenStudy (anonymous):

-5 ?

OpenStudy (radar):

Did you want to use the quadratic formula? If so you need to have an equation, something with an equal sign, What you have written is an expression, it can be factored, but there is no zeros unless itself is equal to zero.

OpenStudy (anonymous):

I think I already have one factored enough. Can you help me with the second one?

OpenStudy (radar):

*I only saw one. The first one factored to (x+5)(2x+3)

OpenStudy (radar):

What does the second one look like?

OpenStudy (radar):

@unheard do you have a second one?

OpenStudy (radar):

Good luck with your studies.

OpenStudy (anonymous):

@jim_thompson5910 can you help me with 5x^2-2x-7 ?

OpenStudy (espex):

What are the factors of \(5x^2\)? What are the factors of 7?

jimthompson5910 (jim_thompson5910):

another way to factor is to first solve 5x^2-2x-7 = 0 for x (use the quadratic formula) once you have the roots, you can use them to factor

OpenStudy (anonymous):

The factors of 7 are 1 and 7 and the factors of 5 are 1 and 5. I don't know about the x^2 part >.<

OpenStudy (anonymous):

@eSpeX ?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what do you get when you solve 5x^2-2x-7 = 0 for x

OpenStudy (anonymous):

Can you break the problem up so it makes it easier for me to work it out. > . <

jimthompson5910 (jim_thompson5910):

are you familiar with the quadratic formula?

OpenStudy (anonymous):

Kinda. It's really hard.

jimthompson5910 (jim_thompson5910):

5x^2-2x-7 = 0 a = 5 b = -2 c = -7

jimthompson5910 (jim_thompson5910):

the discriminant is D = b^2 - 4ac D = (-2)^2 - 4(5)(-7) D = ???

OpenStudy (anonymous):

4 - (-140)

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

144

jimthompson5910 (jim_thompson5910):

since that is a perfect square, this means that you'll get 2 rational solutions this means the original expression can be factored

jimthompson5910 (jim_thompson5910):

if you didn't get a perfect square, then it couldn't be factored and you can move on and not bother to try to factor it

jimthompson5910 (jim_thompson5910):

anyways, you then take the square root of this to get sqrt(144) = 12

jimthompson5910 (jim_thompson5910):

you will then have x = (-b+sqrt(D))/(2a) x = (-(-2)+sqrt(144))/(2*5) x = (2+12)/(10) x = (14)/(10) x = 7/5 that's one solution

jimthompson5910 (jim_thompson5910):

the other solution is x = (-b-sqrt(D))/(2a) x = (-(-2)-sqrt(144))/(2*5) x = (2-12)/(10) x = (-10)/(10) x = -1

OpenStudy (johnweldon1993):

Thanks for taking it over @jim_thompson5910 !

jimthompson5910 (jim_thompson5910):

now that you know the two solutions are x = 7/5 and x = -1, we can do this x = 7/5 or x = -1 5x = 7 or x = -1 5x - 7 = 0 or x +1 = 0 (5x - 7)(x + 1) = 0

jimthompson5910 (jim_thompson5910):

so this shows that 5x^2-2x-7 = 0 turns into (5x - 7)(x + 1) = 0 which means 5x^2-2x-7 factors to (5x - 7)(x + 1)

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

yw